OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x3 – 2x2 – 16x + 32 . x = –4, x = –2, x = 2 x = –2, x = 2, x = 4 x = –4, x = –2, x = 4 x = –4, x = 2, x = 4

sam (.sam.):

x3 – 2x2 – 16x + 32=0 factor (x-4) (x^2+2x-8)=0 (x-4) (x-2) (x+4)=0 x=4 x=2 x=-4